OpenStudy (anonymous):

What is the slope of the line which passes through (4, 5) and (0, 1)?

OpenStudy (anonymous):

The slope is given by the equations \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

\[SLOPE OF A LINE (M) = Y2-Y1\div X2-X1\] --( Where (X1, Y1) & (X2,Y2) are the given coordinates) For your given problem : (X1, Y1) is (4,5) & (Y1,Y2) is (0,1) ** You can also choose vice versa if you want. That shouldnt be a problem. So your Slope (M) as calculated by the above given formula should be M = (5-1) / (4-0) = 4/4 =1 Answer: Slope (M) = 1